### v1

@@ 5.1 Confidence Sets and Maximal Matrix
For any confidence set $\mathcal{C}^{t}(\prec,p,H_{t})$, there exists a unique element-wise maximal matrix $P^{t}\in\mathcal{C}^{t}$ such that $P^{t}_{u,k}\geq P_{u,k}$ for all $P\in\mathcal{C}^{t}$ and all $u\in[U],k\in[K]$. Furthermore, this matrix is given by
$P^{t}_{u,k}=\begin{cases}\min_{i:k\preceq i\preceq p(u)}UCB_{u,i}(t),&k\preceq p(u)\\ \min_{i:p(u)\preceq i\preceq k}UCB_{u,i}(t),&p(u)\preceq k\end{cases}.$
We refer to the above matrix as the maximal matrix of $\mathcal{C}^{t}$. This maximal matrix could not be exploited in the previous section, as its characterization fundamentally relies on knowledge of both the order $\prec$ and the peaks $p(\cdot)$.
5.2 The Match-via-Maximal Algorithm
Algorithm 3 Match-via-Maximal (MvM)
1:Order $\prec$, peaks $p^{\star}(\cdot)$
2:for $t=1$ to $T$ do
3:  Construct the maximal matrix $P^{t}$
4:  Select $\pi_{t}=$ SP-Matching($P^{t}$)
5:  Observe rewards and update history $H_{t}$
Next, we present a combinatorial UCB-based algorithm, which we refer to as MvM (Algorithm 3). MvM acts optimistically—in each round, it selects the optimal matching w.r.t. the corresponding maximal matrix. The regret analysis of MvM involves standard UCB-like arguments. By the construction of the confidence sets and applying Hoeffding’s inequality with a union bound over all $(u,k)$ pairs and time steps, the actual expected reward matrix $\Theta$ is in $\mathcal{C}^{t}$ for all $t\in[T]$ with probability at least $1-\frac{2KU}{T^{3}}$. Employing the clean event technique, we ignore the complementary (bad) event, which occurs with low probability. In every round $t$, we have
$\max_{\pi\in\Pi}V(\pi;\Theta)-V(\pi_{t};\Theta)\leq V(\pi_{t};P^{t})-V(\pi_{t};\Theta)\leq\sum_{u\in[U]}2\sqrt{\frac{2\ln T}{n_{u,\pi_{t}(u)}}}.$ (5)
Summing Inequality (5) over all rounds, we obtain the following theorem:
Theorem 8.
For any SP instance with known SP order $\prec$ and peaks $p(\cdot)$, MvM achieves regret at most $O(U\sqrt{TK\ln T})$ with per-round runtime of $O(K^{2}U+K^{2}B)$.
Unknown separated peaks.
The assumption of known peaks can be relaxed when each user’s peak is separated from their second-best option by a positive gap $\Delta_{\text{peak}}$. We stress that $\Delta_{\text{peak}}$ is unknown to the algorithm. Under this condition, we can utilize a mixed exploration-UCB approach to obtain a regret of $O\left(\frac{UK\ln T}{\Delta_{\text{peak}}^{2}}+U\sqrt{TK\ln T}\right)$, where the first term represents the peak-identification cost and the second term corresponds to applying Algorithm 3 for the remaining rounds. We defer the full details to Appendix C.
6 Lower Bounds for Known Order
In this section, we show that the class of SP instances still poses a significant challenge for the learning algorithm. We demonstrate this by establishing lower bounds on the regret for the most lenient information knowledge: when the SP order is known. W.l.o.g., we assume the identity SP order and formulate our results in terms of PSP instances.
We present our results in the following theorem, which provides a lower bound on the worst-case expected regret for both known and unknown peaks scenarios.
Theorem 9.
These lower bounds naturally apply to the unknown order case as well, since algorithms operate with strictly less information in that setting. Notably, the lower bound for unknown user peaks matches that of the general preferences case, as detailed in Appendix E.
7 Discussion
We studied a budgeted matching problem under single-peaked preferences. While the general offline problem is NP-hard to approximate within a constant factor, imposing a single-peaked structure enables polynomial-time optimization. When the SP structure is unknown, we introduce an explore-then-commit approach that first extracts an approximate order, then commits to a near-optimal policy, yielding $\tilde{O}(UKT^{2/3})$ regret. For the more lenient case of known SP order and user peaks, we developed an optimistic algorithm based on maximal matrices that achieves $\tilde{O}(U\sqrt{TK})$ regret. Furthermore, we developed lower bounds for the case of known and unknown user peaks. We also conducted an experimental validation, which is deferred to Appendix F due to space constraints.
We identify several interesting directions for future work. First, comparing our algorithmic results and the lower bounds, we observe several gaps. For the unknown SP order, EMC achieves a regret of $\tilde{O}(UKT^{2/3})$, while our lower bound is $\Omega(U\sqrt{TK})$. Future work could either propose an efficient algorithm with better regret or introduce a stronger lower bound under efficient implementation. For instance, solving the optimization problem in Equation 4 for confidence sets that do not require the SP order (and are thus non-convex) would lead to the same $\tilde{O}(U\sqrt{TK})$ regret. For known order and known peaks, our MvM algorithm achieves $\tilde{O}(U\sqrt{TK})$ regret (Theorem 8), which exceeds the lower bound by a factor of $\sqrt{\ln T}\cdot\max\{\sqrt{K},U\}$. Future work could also resolve this gap.
Second, our results suggest that single-peaked preferences could potentially simplify other computationally challenging problems in online learning, suggesting an alternative route to a solution rather than $\alpha$-regret. Finally, previous work has explored generalizations of single-peaked preferences to more complex structures, including single-peaked preferences on trees [32] and cycles [29]. Extending our algorithmic framework and theoretical guarantees to these richer preference structures presents an intriguing challenge.
Appendix A Deferred Proofs for Section 3
See 1
Proof of Lemma 1.
@@ Appendix A Deferred Proofs for Section 3
Remark 1.
Lemma 1 is robust to the arbitrary choice of the peak when a user has several maximizers. Suppose $u$ has two peak indices $x<y$ with $\theta_{u,x}=\theta_{u,y}=\max_{k\in[K]}\theta_{u,k}$, and fix a selected set $S=\{k_{1}<k_{2}<\cdots<k_{m}\}$. Then the value of the best matching of $u$ against $S$ is independent of whether we instantiate $p(u)=x$ or $p(u)=y$. Specifically, exactly one of the following holds:
  • There exists $j\in\{1,\ldots,m-1\}$ with $k_{j}\leq x\leq y\leq k_{j+1}$ and no $k_{\ell}\in(x,y)$. Then by unimodality, any arm outside $\{k_{j},k_{j+1}\}$ is dominated by one of them, so $\pi^{\star}(u)\in\{k_{j},k_{j+1}\}$, regardless of whether $p(u)=x$ or $p(u)=y$.
  • $y\leq k_{1}$. Values are nonincreasing to the right of the peaks, hence $\pi^{\star}(u)=k_{1}$, regardless of the tie-break.
  • $k_{m}\leq x$. Values are nondecreasing to the left of the peaks, hence $\pi^{\star}(u)=k_{m}$, regardless of the tie-break.
  • There exist indices $i\leq j$ with $x\leq k_{i}\leq\cdots\leq k_{j}\leq y$. Since all $k\in[x,y]$ attain the peak value from the SP property, any $k\in\{k_{i},\ldots,k_{j}\}$ is optimal, regardless of the tie-break. The lemma asserts that $\pi^{\star}(u)\in\{k_{i-1},k_{i}\}$ if the peak is set to $x$, and $\pi^{\star}(u)\in\{k_{j},k_{j+1}\}$ if the peak is set to $y$. In either case, the attained value $\max_{k\in S}\theta_{u,k}$ remains unchanged.
In every case, $u$’s attained value is invariant to the choice of the peak.
@@
By Lemma 1, users with peaks in the interval $(i,k]$ are optimally assigned to either arm $i$ or arm $k$, contributing exactly $G_{i,k}$ to the total reward. Users with peaks in $\{0,1,\ldots,i\}$ contribute optimally according to $\text{OPT}(i,b-c_{k})$ by definition. Since the fictive arm 0 has zero cost, the constraint $b\geq c_{i}+c_{k}$ is always satisfiable with $i=0$. Therefore,
@@
which accounts for users whose peaks exceed the rightmost selected arm due to the second case in Lemma 1. This ensures that the algorithm correctly computes the maximum reward for the given budget $B$. The backtracking reconstructs the optimal matching, from which we can exclude arm 0 if it was somehow included by picking any other arm for that user.
@@
Remark 2.
SP-Matching is invariant to the arbitrary choice of the peak when a user has several maximizers as well. More specifically, notice that every choice of the last arm $k$ in the subset (Line (9)) corresponds to a subset of selected arms, and a matching that assigns each user based on Lemma 1. Thus, according to Remark 1, the attained value of each user in this subset is independent of the choice of the peak. The choice of the peak only affects when we account for this user’s contribution to the total value, not the contribution itself. Therefore, running the algorithm with different valid peak indices yields the same final value $V^{\star}$.
Appendix B Deferred Proofs for Section 4
See 3
Proof of Proposition 3.
We prove each direction separately.
SP $\Rightarrow$ 0-ASP: Let $P$ be an SP matrix with respect to an order $\prec$, user $u\in[U]$, and a triplet of arms $i\prec j\prec\ell$. We divide the analysis into cases based on the relative position of $j$ w.r.t. the user’s peak $p(u)$, leveraging the unimodality of the preference vector $P_{u,(\cdot)}$. If $i\prec j\preceq p(u)$, then $P_{u,j}\geq P_{u,i}$; additionally, if $p(u)\preceq j\prec\ell$, then $P_{u,j}\geq P_{u,\ell}$. In either case, we have
$P_{u,j}\geq\min\{P_{u,i},P_{u,\ell}\},$
which implies that $P$ is 0-ASP with respect to $\prec$.
0-ASP $\Rightarrow$ SP: Let $P$ be a 0-ASP matrix with respect to an order $\prec$, and user $u\in[U]$. Denote $p(u)=\operatorname*{arg\,max}_{k\in[K]}P_{u,k}$. Then, for every arms $i,j$ such that $i\prec j\prec p(u)$, we have
$P_{u,j}\geq\min\{P_{u,i},P_{u,p(u)}\}=P_{u,i},$
Which implies that $P_{u,i}\leq P_{u,j}\leq P_{u,p(u)}$. Similarly, for every arms $i,j$ such that $p(u)\prec i\prec j$ it holds that $P_{u,p(u)}\geq P_{u,i}\geq P_{u,j}$. Thus, $P_{u,(\cdot)}$ is unimodal with peak at $p(u)$, and therefore $P$ is SP with respect to $\prec$. ∎
See 4
Proof of Lemma 4.
@@ Appendix B Deferred Proofs for Section 4
It is straightforward to verify that $P$ is SP w.r.t. $\prec$, as $P_{u,(\cdot)}$ is non-decreasing up to $p(u)$ and non-increasing afterward. Now, focus on an entry $(u,k)$ of $P$ such that $k\prec p(u)$. Then, by the construction of $P$, we have:
$P_{u,k}=\max_{i\in[K]:i\preceq k}\tilde{P}_{u,i}\leq\max\{\tilde{P}_{u,k},\max_{i\in[K]:i\prec k}\tilde{P}_{u,k}+\delta\}=\tilde{P}_{u,k}+\delta,$
where the inequality follows from employing Inequality ((2)) using $i\prec k\prec p(u)$ for every $i\prec k$. A similar argument holds for entries $(u,k)$ such that $p(u)\prec k$. Therefore, for every pair $(u,k)$ it holds that $P_{u,k}-\tilde{P}_{u,k}\leq\delta$. Additionally, from construction, we always have $P_{u,p(u)}\geq\tilde{P}_{u,p(u)}$. Thus, we can conclude that $\lvert P_{u,k}-\tilde{P}_{u,k}\rvert\leq\delta$ for every $u\in U$ and $k\in[K]$. ∎
See 5
Proof of Lemma 5.
Let $\prec_{P}$ be an SP order of $P$. Recall that Proposition 3 implies that $P$ is 0-ASP with respect to $\prec_{P}$. We will show that $\prec_{P}$ satisfies all the constraints that can be imposed in Line (7) of the Extract-Order algorithm. Fix a user $u$, and recall that we denote the sequence of $u$’s sorted preferences as $\tilde{P}_{u,k_{1}^{u}},\tilde{P}_{u,k_{2}^{u}},\ldots,\tilde{P}_{u,k_{K}^{u}}$. Let $m$ be an index such that $\tilde{P}_{u,k_{m}^{u}}>\tilde{P}_{u,k_{m+1}^{u}}+2\varepsilon$. Because $\lVert\tilde{P}-P\rVert_{\infty}\leq\varepsilon$, for all $i\in[m]$ and $j\in[m+1,K]$ it holds that
$P_{u,k_{i}^{u}}-P_{u,k_{j}^{u}}\geq\tilde{P}_{u,k_{i}^{u}}-\tilde{P}_{u,k_{j}^{u}}-2\varepsilon>0.$ (6)
Since Proposition 3 ensures that $P$ is 0-ASP with respect to $\prec_{P}$, if any arm $k_{j}^{u}$ with $j\in[m+1,K]$ were to appear between two arms $k_{i}^{u},k_{\ell}^{u}$ with $i,\ell\in[m]$ in the order $\prec_{P}$, then the 0-ASP condition would require
$P_{u,k_{j}^{u}}\geq\min\{P_{u,k_{i}^{u}},P_{u,k_{\ell}^{u}}\}.$
However, this violates (6). Therefore, the sequence $k_{1}^{u},k_{2}^{u},\ldots,k_{m}^{u}$ must appear contiguously in $\prec_{P}$. Thus, every constraint that might be imposed in Line (7) is naturally satisfied by $\prec$.
After showing that a feasible order exists, we show that $\tilde{P}$ is $(2K\varepsilon)$-ASP w.r.t. any output order $\prec$. Fix any user $u$ and arms $i\prec j\prec\ell$. We need to show that
$\tilde{P}_{u,j}\geq\min\{\tilde{P}_{u,i},\tilde{P}_{u,\ell}\}-2K\varepsilon.$
Since $i\prec j\prec\ell$, the algorithm couldn’t have imposed a contiguity constraint that contains both $i$ and $\ell$ but excludes $j$. This means that in user $u$’s sorted preference sequence $\tilde{P}_{u,k_{1}^{u}}\geq\cdots\geq\tilde{P}_{u,k_{K}^{u}}$, there is no index $m$ such that $\{k_{1}^{u},\ldots,k_{m}^{u}\}$ contains $\{i,\ell\}\setminus\{j\}$ and $\tilde{P}_{u,k_{m}^{u}}-\tilde{P}_{u,k_{m+1}^{u}}>2\varepsilon$.
Without loss of generality, assume $\tilde{P}_{u,i}\leq\tilde{P}_{u,\ell}$. The difference $\tilde{P}_{u,j}-\tilde{P}_{u,i}$ is bounded by the sum of consecutive gaps in the sorted sequence between the positions of $j$ and $i$. Since no such gap exceeds $2\varepsilon$, and there are at most K1K-1 gaps, we have:
$\tilde{P}_{u,i}-\tilde{P}_{u,j}\leq 2K\varepsilon,$
which completes the proof. ∎
See 6
Proof of Theorem 6.
Using Hoeffding’s inequality and taking a union bound over all pairs $(u,k)$, we can ensure that for all $u\in[U],k\in[K]$,
$\lvert\bar{\theta}_{u,k}-\theta_{u,k}\rvert\leq\sqrt{\frac{2\ln T}{N}}\quad$ (7)
holds with probability at least $1-\frac{2KU}{T^{4}}$. As the complementary event contributes at most $TU$ to the regret, its effect on the regret is negligible.
Let $\tilde{\Theta}$ be the SP matrix constructed in Line (4). Lemma 4 and Lemma 5 ensure that
$\lVert\tilde{\Theta}-\bar{\Theta}\rVert_{\infty}\leq 2K\sqrt{\frac{2\ln T}{N}}.$ (8)
Combining (7) and (8), we obtain:
$\lVert\tilde{\Theta}-\Theta\rVert_{\infty}\leq(2K+1)\sqrt{\frac{2\ln T}{N}}.$ (9)
@@
Now, let $\tilde{\pi}$ be the matching computed in Line (5), and let $\pi^{\star}$ be the optimal matching w.r.t. $\Theta$. Using (9), we can bound the regret of a single commitment round ($t>KN$) as follows:
$V(\pi^{\star};\Theta)-V(\tilde{\pi};\Theta)\leq V(\pi^{\star};\tilde{\Theta})-V(\tilde{\pi};\tilde{\Theta})+2U(2K+1)\sqrt{\frac{2\ln T}{N}}\leq 2U(2K+1)\sqrt{\frac{2\ln T}{N}},$
@@
$R_{T}\leq O\left(KNU+T\cdot 2U(2K+1)\sqrt{\frac{2\ln T}{N}}\right).$

### v2

@@ A.4 Efficient Approximation via Alpha-Regret
The reformulated reward function
$f(S)=\sum_{u\in[U]}\max_{k\in S}\Theta_{u,k}$
is submodular and monotone.
Proof of Lemma˜16.
We prove both properties separately.
$f(M\cup\{k\})-f(M)\geq f(N\cup\{k\})-f(N).$
$f(M\cup\{k\})-f(M)=\sum_{u\in[U]}\max\{0,\Theta_{u,k}-\theta^{M}_{u}\},$
$f(N\cup\{k\})-f(N)=\sum_{u\in[U]}\max\{0,\Theta_{u,k}-\theta^{N}_{u}\}.$
For each user $u$, since $\theta^{M}_{u}\leq\theta^{N}_{u}$:
Summing over all users gives the desired inequality.
Proposition 17.
Proof of Proposition˜17.
Regret analysis. We reformulate our problem as submodular maximization over subsets, where we select $S\subseteq[K]$ subject to the budget constraint $\sum_{k\in S}c_{k}\leq B$ and each user receives their most preferred item from the selected subset. The reformulated reward function $f(S)=\sum_{u\in[U]}\max_{k\in S}\Theta_{u,k}$ is submodular and monotone by Lemma˜16. Under knapsack constraints, the Greedy+Max algorithm of Yaroslavtsev et al. [37] achieves a $\frac{1}{2}$-approximation and is $\left(\frac{1}{2},\frac{1}{2}+\tilde{K}+2\beta\right)$-robust in the sense of Nie et al. [30], where $\tilde{K}=\min\{K,B/c_{\min}\}$ and $\beta=B/c_{\min}$. Applying the C-ETC framework of Nie et al. [30] with normalized rewards $\tilde{f}(S)=\frac{1}{U}f(S)$ yields the bound $O(U\cdot(\frac{1}{2}+\tilde{K}+2\beta)^{2/3}(\tilde{K}K)^{1/3}T^{2/3}\log(T)^{1/3})$. To obtain our clean bound, we use $(\frac{1}{2}+\tilde{K}+2\beta)^{2/3}\leq(K+\beta)^{2/3}$ and $(\tilde{K}K)^{1/3}\leq K^{2/3}$, then apply $(K+\beta)^{2/3}\cdot K^{2/3}\leq(K+\beta)^{4/3}$ to get the stated bound.
Appendix B Proof of Theorem 3
See 3
Proof of Theorem˜3.
We prove the lower bounds by constructing specific hard instances for each regime.
Case 1: Known Order and Peaks ($R_{T}=\Omega(U\sqrt{T})$).
We construct a family of hard instances parameterized by $\omega=\big(({\mu}_{i},{\nu}_{i})\big)_{i=1}^{U}\in[1/4,3/4]^{2U}$. To illustrate the construction, consider a "base case" consisting of three users and four arms with a budget $B=3$ and costs $c=(1,3,1,1)$. The reward structure for this case is shown in Table˜1.
Table 1: Reward structure for the base case (Case 1). Costs are shown in parentheses.
Arm 1 ($c_{1}=1$) Arm 2 ($c_{2}=3$) Arm 3 ($c_{3}=1$) Arm 4 ($c_{4}=1$)
User 1 $\mu$ 1 $\nu$ 0
User 2 0 0 0 1
User 3 0 0 0 1
The full instance comprises $U$ copies of User 1 (indexed $1,\dots,U$) and $2U$ copies of Users 2 and 3 (indexed $U+1,\dots,3U$). The instance is PSP with order $1\prec 2\prec 3\prec 4$. Users $1\dots U$ peak at Arm 2, while the others peak at Arm 4. Due to the budget $B=3$, any feasible matching must either select Arm 2 alone (cost 3), or a subset of $\{1,3,4\}$. The optimal matching selects $\{1,3,4\}$, assigning users $1\dots U$ to $\operatorname*{arg\,max}(\mu_{i},\nu_{i})$ and the rest to Arm 4. The optimal expected reward is $V^{\star}(\omega)=\sum_{i=1}^{U}\max\{\mu_{i},\nu_{i}\}+2U$.
Consider an algorithm $\mathcal{A}$ generating matchings $\pi^{t}$. Let $T_{2}=\{t\in[T]:\text{Im}(\pi^{t})=\{2\}\}$ be the rounds where Arm 2 is selected, and $T_{1}=[T]\setminus T_{2}$. In rounds $t\in T_{2}$, the reward is exactly $U$. In rounds $t\in T_{1}$, users $U+1\dots 3U$ contribute at most $2U$ (if they are being matched to Arm 4), while user $i\in[U]$ contributes based on their assignment to 1,3,4. Decomposing the cumulative reward, similar to the classical bandit analysis, we can upper-bound the reward by
$\mathbb{E}_{\omega}^{\mathcal{A}}\left[\sum_{t=1}^{T}V(\pi^{t};\Theta^{\omega})\right]\leq\mathbb{E}_{\omega}^{\mathcal{A}}\left[U\lvert T_{2}\rvert+2U\lvert T_{1}\rvert+\sum_{i=1}^{U}\left(\mu_{i}N_{1}^{i}+\nu_{i}N_{3}^{i}\right)\right],$
where $N_{j}^{i}$ is the number of times user $i$ is matched to arm $j$. The regret is then bounded by:
$R_{T}\geq\mathbb{E}_{\omega}^{\mathcal{A}}\left[U\lvert T_{2}\rvert+\lvert T_{1}\rvert\sum_{i=1}^{U}\max\{{\mu}_{i},{\nu}_{i}\}-\sum_{i=1}^{U}\left(\mu_{i}N_{1}^{i}+\nu_{i}N_{3}^{i}\right)\right].$
If $\mathbb{E}[|T_{2}|]\geq\sqrt{T}$, the first term yields $\Omega(U\sqrt{T})$. If $\mathbb{E}[|T_{2}|]<\sqrt{T}$, then $\mathbb{E}[|T_{1}|]=\Omega(T)$, and the remaining terms correspond to the regret of $U$ independent 2-armed bandit problems played for $\Omega(T)$ rounds. By the standard minimax lower bound [2], this scales as $\Omega(U\sqrt{T})$.
Case 2: Known Order and Peaks ($R_{T}=\Omega(\sqrt{TK})$).
To capture the dependency on $K$, assume $U=2$, budget $B=1$ (forcing a choice of exactly one arm per round), and costs $c_{k}=1$. We construct a family of hard PSP instances parameterized by $i^{\star}\in[K]$. For each $i^{\star}$, the reward structure is defined as follows: User 1 has preferences $1/2+\Delta$ for arms $k\leq i^{\star}$ and $1/2$ for arms $k>i^{\star}$, while User 2 has preferences $1/2$ for arms $k<i^{\star}$ and $1/2+\Delta$ for arms $k\geq i^{\star}$.
Each instance is PSP: User 1’s preferences are non-increasing with a peak at arm 1, and User 2’s preferences are non-decreasing with a peak at arm $K$. For any choice of $i^{\star}$, arm $i^{\star}$ is the unique optimal arm with total reward $1+2\Delta$, while all other arms yield $1+\Delta$. Identifying the optimal arm among $K$ possibilities reduces to the classical $K$-armed bandit problem. Following the standard change-of-measure argument via Bretagnolle-Huber (e.g., Lattimore and Szepesvári [25, Chapter 15]) with $\Delta\asymp\sqrt{K/T}$, we obtain a lower bound of $\Omega(\sqrt{TK})$. Note that although each round yields two reward observations rather than one, this only affects the KL-divergence by a constant factor and does not change the order of the lower bound. Combining this with Case 1, we have a regret lower bound of $\Omega(\max\{U\sqrt{T},\sqrt{TK}\})$.
Case 3: Unknown Peaks ($R_{T}=\Omega(U\sqrt{TK})$).
If the peaks are unknown, we rely on the classic construction of the multi-armed bandit lower bound (e.g., Slivkins [35, Chapter 2]). Specifically, we construct an instance where each user $u$ has a unique, unknown peak $p(u)$ with reward $1/2+\Delta$, while all other arms offer reward $1/2$. Observe that such preference profiles are trivially PSP regardless of the underlying order of arms, as they are constant everywhere except at a single point (the peak). Since the distributions are independent across users, learning the peak for one user provides no information about the others. This reduces to $U$ independent $K$-armed bandit instances. By the standard minimax lower bound, each user contributes $\Omega(\sqrt{TK})$ to the regret, resulting in a total regret of $\Omega(U\sqrt{TK})$. ∎
Appendix C Deferred Proofs from Section˜4
See 4
Proof of Lemma˜4.
@@ Appendix C Deferred Proofs from Section˜4
By Lemma˜4, users with peaks in the interval $(i,k]$ are optimally assigned to either arm $i$ or arm $k$, contributing exactly $G_{i,k}$ to the total reward. Users with peaks in $\{0,1,\ldots,i\}$ contribute optimally according to $\text{OPT}(i,b-c_{k})$ by definition. Since the fictive arm 0 has zero cost, the constraint $b\geq c_{i}+c_{k}$ is always satisfiable with $i=0$. Therefore,
@@
which accounts for users whose peaks exceed the rightmost selected arm due to the second case in Lemma˜4. This ensures that the algorithm correctly computes the maximum reward for the given budget $B$. The backtracking reconstructs the optimal matching, from which we can exclude arm 0 if it was somehow included by picking any other arm for that user.
@@
Appendix D Deferred Proofs and Details from Section˜5
See 6
Proof of Lemma˜6.
First, observe that $P^{t}$ is SP by construction (as the values are defined via a running minimum moving away from the peak) and satisfies $P^{t}\leq\textnormal{UCB}$ by definition.
Next, assume towards contradiction that there exists a matrix $P\in\mathcal{C}$ and a pair $(u,k)$ such that $P_{u,k}>P^{t}_{u,k}$. By the definition of $P^{t}$, there exists an index $k^{\prime}$ with either $k\preceq k^{\prime}\preceq p(u)$ or $p(u)\preceq k^{\prime}\preceq k$ such that $P^{t}_{u,k}=\textnormal{UCB}_{u,k^{\prime}}(t)$. From our assumption, we have $P_{u,k}>\textnormal{UCB}_{u,k^{\prime}}$. Since $P\in\mathcal{C}$, Definition˜3 ensures that $P_{u,k^{\prime}}\leq\textnormal{UCB}_{u,k^{\prime}}(t)$. Combining these inequalities yields $P_{u,k}>\textnormal{UCB}_{u,k^{\prime}}(t)\geq P_{u,k^{\prime}}$. However, the SP property of $P$ with peak at $p(u)$ implies $P_{u,k}\leq P_{u,k^{\prime}}$ (since both arms lie on the same side of the peak), contradicting $P_{u,k}>P_{u,k^{\prime}}$.
Finally, if $\mathcal{C}^{t}$ is non-empty, then for any $P\in\mathcal{C}^{t}$, we have $P^{t}\geq P\geq\textnormal{LCB}$, ensuring $P^{t}\in C^{t}$ and completing the proof. ∎
See 7
Proof of Theorem˜7.
The analysis follows the same structure as the UCB-based algorithm in Corollary˜2, with the maximal matrix $P^{t}$ playing the role of the UCB matrix. We condition on the clean event $\mathcal{E}=\{\Theta\in\mathcal{C}^{t}\text{ for all }t\}$, which holds with probability at least $1-\nicefrac{{2UK}}{{T^{3}}}$ by Hoeffding’s inequality and a union bound.
Under $\mathcal{E}$, since $P^{t}$ is element-wise maximal in $\mathcal{C}^{t}$ (Lemma˜6), the per-round regret satisfies Inequality ((3)). Summing over rounds and applying the same regrouping and Jensen’s inequality arguments as in the proof of Corollary˜2 yields $R_{T}=O(U\sqrt{TK\ln T})$.
For the runtime, each round involves updating statistics ($O(U)$), constructing $P^{t}$ via Lemma˜6 ($O(UK)$), and running SP-Matching ($O(K^{2}(U+B))$ by Theorem˜5). The complexity is dominated by the matching step. ∎
D.1 Extension to Partial Structural Knowledge
[…]
Appendix E Deferred Proofs from Section˜6
See 8
Proof of Proposition˜8.
First, assume $P$ is SP w.r.t. $\prec$. By definition, for any user $u$, the values $P_{u,k}$ are non-decreasing up to a peak $p(u)$ and non-increasing thereafter. Consequently, for any triplet $i\prec j\prec\ell$, the middle element $j$ must satisfy $P_{u,j}\geq P_{u,i}$ (if $j\preceq p(u)$) or $P_{u,j}\geq P_{u,\ell}$ (if $j\succeq p(u)$). In either case, $P_{u,j}\geq\min\{P_{u,i},P_{u,\ell}\}$, satisfying the 0-ASP condition.
Conversely, suppose $P$ is 0-ASP w.r.t. $\prec$. For any user $u$, let $p(u)$ be an index maximizing $P_{u,\cdot}$. For any $i\prec j\prec p(u)$, the 0-ASP condition implies $P_{u,j}\geq\min\{P_{u,i},P_{u,p(u)}\}=P_{u,i}$, ensuring non-decreasing values to the left of the peak. By symmetry, values are non-increasing to the right of the peak, implying that $P$ is SP. ∎
See 9
Proof of Lemma˜9.
@@ Appendix E Deferred Proofs from Section˜6
The matrix $P$ is SP w.r.t. $\prec$ by construction: for each user $u$, the row $P_{u,(\cdot)}$ is defined as a running maximum from the endpoints toward $p(u)$, ensuring non-decreasing values for $k\preceq p(u)$ and non-increasing values for $k\succeq p(u)$.
It remains to show that $\lVert P-\tilde{P}\rVert_{\infty}\leq\delta$. Consider an entry $(u,k)$ with $k\prec p(u)$. By construction, $P_{u,k}=\max_{i:i\preceq k}\tilde{P}_{u,i}$. If the maximum is achieved at $k$ itself, then $P_{u,k}=\tilde{P}_{u,k}$. Otherwise, the maximum is achieved at some $i\prec k$, and we have $P_{u,k}=\tilde{P}_{u,i}\leq\tilde{P}_{u,k}+\delta$, where the inequality follows from applying Definition˜4 to the triplet $i\prec k\prec p(u)$. A symmetric argument applies when $p(u)\prec k$. Since $P_{u,k}\geq\tilde{P}_{u,k}$ by construction (as the maximum includes $\tilde{P}_{u,k}$ itself), we conclude $|P_{u,k}-\tilde{P}_{u,k}|\leq\delta$ for all $(u,k)$. ∎
See 10
Proof of Lemma˜10.
We first establish the existence of an order that satisfies all the constraints; we do so via the SP order of $P$, which we denote as $\prec_{P}$. Fix some $u\in U$ for which a contiguity constraint on the set of arms $S=\{k_{1}^{u},\ldots,k_{m}^{u}\}$ was imposed. Due to Lines 56, we must have $\min_{k\in S}\tilde{P}_{u,k}-\max_{k^{\prime}\notin S}\tilde{P}_{u,k^{\prime}}>2\varepsilon$. Since $\lVert\tilde{P}-P\rVert_{\infty}\leq\varepsilon$, this implies $\min_{k\in S}P_{u,k}-\max_{k^{\prime}\notin S}P_{u,k^{\prime}}>0$. Thus, because $P$ is 0-ASP w.r.t. $\prec_{P}$, $\prec_{P}$ must satisfy this constraint as well; otherwise, as was previously discussed, we could have found a triplet of arms that contradicts Definition˜4.
Next, let $\prec$ be a returned order (one must exist, as our first step suggests). Fix $u$ and arms $i\prec j\prec\ell$. We must show $\tilde{P}_{u,j}\geq\min\{\tilde{P}_{u,i},\tilde{P}_{u,\ell}\}-2K\varepsilon$. Assume w.l.o.g. that $\tilde{P}_{u,i}\leq\tilde{P}_{u,\ell}$. Consider the arms sorted by preference $k_{1}^{u},\dots,k_{K}^{u}$. If there were any index $m$ such that $\{k_{1}^{u},\dots,k_{m}^{u}\}$ includes $i$ and $\ell$ but excludes $j$, and the gap $\tilde{P}_{u,k_{m}^{u}}-\tilde{P}_{u,k_{m+1}^{u}}>2\varepsilon$, the algorithm would constrain this set to be contiguous. This would force $j$ outside the interval between $i$ and $\ell$, contradicting $i\prec j\prec\ell$. Therefore, no gap exceeding $2\varepsilon$ exists in the sorted sequence between the values $\tilde{P}_{u,i}$ and $\tilde{P}_{u,j}$. Summing these gaps (at most $K$) yields $\tilde{P}_{u,i}-\tilde{P}_{u,j}\leq 2K\varepsilon$. ∎
See 11
Proof of Theorem˜11.
Let $\varepsilon=\sqrt{2\ln T/N}$. By Hoeffding’s inequality and a union bound, the clean event $\mathcal{E}=\{\lVert\bar{\Theta}-\Theta\rVert_{\infty}\leq\varepsilon\}$ holds with probability at least $1-\nicefrac{{2KU}}{{T^{4}}}$. Its complement, $\mathcal{E}^{C}$, occurs with low probability and contributes at most $TU$ to the overall regret; thus, we condition the rest of the proof on $\mathcal{E}$. Under $\mathcal{E}$, the approximation guarantees from Lemmas˜9 and 10 combined imply that the constructed matrix $\tilde{\Theta}$ satisfies $\lVert\tilde{\Theta}-\bar{\Theta}\rVert_{\infty}\leq 2K\varepsilon$. By the triangle inequality, the total estimation error is bounded by
$\lVert\tilde{\Theta}-\Theta\rVert_{\infty}\leq(2K+1)\sqrt{\nicefrac{{2\ln T}}{{N}}}.$ (6)
@@
Now, let $\tilde{\pi}$ be the matching computed in Line (5), and let $\pi^{\star}$ be the optimal matching w.r.t. $\Theta$. Using ˜(6), we can bound the regret of a single commitment round ($t>KN$) as follows:
$V(\pi^{\star};\Theta)-V(\tilde{\pi};\Theta)\leq V(\pi^{\star};\tilde{\Theta})-V(\tilde{\pi};\tilde{\Theta})+2U(2K+1)\sqrt{\nicefrac{{2\ln T}}{{N}}}\leq 2U(2K+1)\sqrt{\nicefrac{{2\ln T}}{{N}}},$
@@
$R_{T}\leq O\left(KNU+T\cdot 2U(2K+1)\sqrt{\nicefrac{{2\ln T}}{{N}}}\right).$